# Streamlining Redundant Layers to Compress Large Language Models

### v3

@@ Abstract
This paper introduces LLM-Streamline, a novel layer pruning approach for large language models. It is based on the observation that different layers have varying impacts on hidden states, enabling the identification of less important layers. LLM-Streamline comprises two parts: layer pruning, which removes consecutive layers with the lowest importance based on target sparsity, and layer replacement, where a lightweight network is trained to replace the pruned layers to mitigate performance loss. Additionally, a new metric called “stability” is proposed to address the limitations of accuracy in evaluating model compression. Experiments show that LLM-Streamline surpasses previous state-of-the-art pruning methods in both accuracy and stability.
Figure 1: The left figure illustrates the LLM-Streamline workflow, which includes layer pruning to remove consecutive layers and layer replacement where a lightweight network is trained to replace the pruned layers. The right figure presents the comparison results of LLM-Streamline with previous state-of-the-art (SOTA) pruning methods on 12 classification benchmarks (details in Section 4.2) by pruning 25% of the parameters on Llama2-7B. LLM-Streamline achieves 11.2% higher relative accuracy, which is the accuracy of the compressed model compared to the original model, than the previous SOTA.
@@ 1 Introduction
Large language models (LLMs) based on transformer architecturesvaswani2017attention have garnered significant attention and are employed across various domains and tasks. However, as LLMs increase in size, their hardware requirements escalate substantially, thereby constraining their applicability and impeding their deployment in real-world scenarios.
@@
To reduce the hardware requirements for deploying LLMs, research efforts focus on developing compact models that maintain high performance, often through model compression techniqueszhu2023survey ; wang2024survey . These compact models significantly reduce memory consumption, computation time, and deployment costs. Current model compression techniques are broadly categorized into knowledge distillationhinton2015distilling ; gou2021knowledge ; li2022apidistill ; huang2022distill ; ho2022distill , model quantizationliu2021quan ; gholami2022survey ; dettmers2022int8 , and model pruninglouizos2017l0 ; chen2023lorashear ; frantar2023sparsegpt ; das2023beyond ; sun2023simple ; xia2023sheared . Quantization compresses the model by quantizing the weights to lower precision. Distillation achieves compression by transferring the capabilities of a larger teacher model to a smaller, more efficient student model. On the other hand, pruning compresses the model by eliminating unimportant parameters and modules.
@@
Current pruning methods for LLMs primarily focus on pruning dense matricesashkboos2024slicegpt , attention headsmichel2019sixteen ; voita2019mha , filtersmccarley2019filter ; prasanna2020filter , and hidden dimensionsxia2023sheared ; van2023llm , with less emphasis on layer pruning. These methods often result in structural irregularities, making the pruned models difficult to store and deploy. In contrast, the layer pruning method, simply modifies the depth of the LLMs***The layers of LLMs are stored in data structures like nn.ModuleList, so layer pruning only requires removing elements from this list, making it easy to implement., is easier to implement. Therefore, exploring an effective layer-wise pruning method is crucial.
The effect of each layer in LLM can be viewed as a transformation of the input hidden state. We investigate the significance of different layers and the potential for pruning based on these transformations. As shown in Fig. 2, we observe that certain contiguous layers have minimal impact on the hidden state, indicating they are less important. This allows for efficient layer pruning by replacing these less important layers with lighter networks.
Based on the empirical observations described above, we propose LLM-Streamline, a straightforward and effective layer-wise pruning algorithm for LLMs. Our algorithm consists of two parts: layer pruning, in which we remove several consecutive layers with the lowest importance from the original model according to the target sparsity; and layer replacement, in which we train a lightweight network to replace the pruned layers to recover the performance degradation due to pruning. We experiment with various structures as lighter networks to replace the pruned layers, including feed-forward neural network (FFN), SwiGLU-based feed-forward neural network (SwiGLU), and transformer layer (Layer). Additionally, we find that removing layers without replacement significantly decreases model performance, highlighting the importance of layer replacement.
@@
Moreover, we observe that the existing accuracy metric for evaluating model compression methods has limitations. For natural language understanding tasks involving multiple-choice classification, the compressed model might guess the correct answers for samples which the original model is uncertain about, leading to an overestimation of model compression performance. To address this, we propose a new metric called “stability” to emphasize the original model’s confidence and the consistency between the model before and after pruning.
@@
  • We propose LLM-Streamline, a layer-wise pruning algorithm that removes less important layers by analyzing their transformations to hidden states in LLMs. To mitigate the potential performance loss caused by pruning, we innovatively employ a lightweight network to approximate the functionality of the pruned layers, thereby compensating for the loss incurred.
  • We propose a new metric called “stability”, which considers both the confidence of the original model in its answers and the consistency between the pruned and original model’s answers. The stability metric more accurately reflects the performance of the pruned model in classification tasks compared to the existing accuracy metric.
  • We conduct experiments on 12 well-known classification benchmarks and 3 generation benchmarks for LLMs. Our results show that for an LLM with 7B parameters and a 25% pruning rate, we can maintain 93% performance in classification tasks and 77% in generation tasks with only 30k data samples, outperforming existing SOTA pruning methods.
Note: As we were finalizing this work, some preprints men2024shortgpt ; gromov2024unreasonable (ShortGPT and UIDL) were posted, presenting a similar pruning method to ours. However, we have also proposed an additional layer replacement component.
Figure 2: The cosine similarity between the input and output hidden states of the layers in OPT-1.3B, OPT-2.7B, OPT-6.7B, and Llama2-7B.
@@ 2 LLM-Streamline
The workflow of the LLM-Streamline framework, shown in Fig. 1 (a), comprises two main steps: layer pruning and layer replacement. First, we prune redundant layers from the LLMs. Then, we train a lightweight network to replace the pruned layers and restore the model’s performance.
@@ 2.1 Layer Redundancy in LLMs
LLMs primarily utilize a transformer architecture, consisting of a series of transformer decoder layers. These layers adopt a residual structure, meaning the effect of each transformer layer can be viewed as a transformation of the input hidden state. Assuming that for the $\ell$-th layer $f$, the parameters are $\theta^{(\ell)}$ and the input hidden states are $\boldsymbol{x}^{(\ell)}$, the effect of layer $f$ can be formalized as
$\boldsymbol{x}^{(\ell+1)}=\boldsymbol{x}^{(\ell)}+f(\boldsymbol{x}^{(\ell)},\theta^{(\ell)}).$ (1)
@@
In (1), the $\ell$-th layer $f$ contributes a transformation $f(\boldsymbol{x}^{(\ell)},\theta^{(\ell)})$ to the input $\boldsymbol{x}^{(\ell)}$. Therefore, we assess the importance of each layer in LLMs by evaluating its impact on the input hidden states. We use the cosine similarity $\cos{(\cdot,\cdot)}$ between inputs $\boldsymbol{x}^{(\ell)}$ and outputs $\boldsymbol{x}^{(\ell+1)}$ as a metric. Essentially, a higher cosine similarity between the inputs and outputs of a layer indicates lower importance, and vice versa. This interpretation arises from the observation that a high cosine similarity suggests the layer’s transformation is minimal, making it more amenable to pruning.
@@
To measure the importance of different layers in LLMs, we randomly select samples from the pre-training data (details in Section 4.1). We then record the hidden states generated by the LLMs for these samples and compute the cosine similarity between the input and output hidden states of each layer. The computation of cosine similarity can be formalized as follows:
$\cos{(\boldsymbol{x}^{(\ell)},\boldsymbol{x}^{(\ell+1)})}=\mathbb{E}_{(\boldsymbol{x}^{(\ell)}_{i},\boldsymbol{x}^{(\ell+1)}_{i})\in\mathcal{D}}\left(\frac{1}{L}\sum_{j=1}^{L}\frac{\boldsymbol{x}^{(\ell)}_{i,j}\cdot\boldsymbol{x}^{(\ell+1)}_{i,j}}{\|\boldsymbol{x}^{(\ell)}_{i,j}\|\cdot\|\boldsymbol{x}^{(\ell+1)}_{i,j}\|}\right),$ (2)
@@
where $\mathcal{D}$ denotes the recorded hidden states from different samples, $\boldsymbol{x}^{(\ell)}_{i}$, $\boldsymbol{x}^{(\ell+1)}_{i}\in R^{d\times L}$ denotes the inputs and outputs hidden states of the $i$-th sample respectively, $d$ denotes the hidden size and $L$ denotes the sequence length of each sample.
@@
To mitigate the effects of model size and model structure, we conduct experiments on four models OPT-1.3B, OPT-2.7B, OPT-6.7Bzhang2022opt and Llama2-7Btouvron2023llama , across two structures and four sizes. The results, illustrated in Fig. 2, show that several consecutive layers in all models exhibit high cosine similarity, indicating a low level of importance.
@@ 2.2 Layer Pruning
As Fig. 2 shows, these least important layers are often contiguous. Then given $n$ determined by the target sparsity, we remove $n$ contiguous layers by finding the initial layer $\ell^{*}(n)$ with the highest cosine similarity for pruning, where $n$ is determined by the target sparsity:
$\ell^{*}(n)=\underset{\ell}{\arg\max}\,\cos(\boldsymbol{x}^{({\ell})},\boldsymbol{x}^{({\ell+n})}),$ (3)
where we randomly select samples from the pre-training data to compute the cosine similarity between $\boldsymbol{x}^{(\ell)}$ and $\boldsymbol{x}^{(\ell+n)}$, as outlined in Eq. (2).
@@ 2.3 Layer Replacement
After the layer pruning process, we aim to replace the pruned layer with a lightweight network that has fewer parameters. The rationale is that these layers contribute only minor transformations to the inputs. Therefore, we hypothesize that the cumulative effect of these layers can be approximated by a lightweight network. Specifically, after identifying the initial layer $\ell^{*}(n)$ for pruning, we use $(\boldsymbol{x}^{(\ell^{*})},\boldsymbol{x}^{(\ell^{*}+n)})$ as the training data to train the lightweight network using mean squared error (MSE) loss, which can be formalized as follows:
@@
Discussion. The potential doubt is why to use layer replacement, instead of using common Parameter-Efficient Fine-Tuning (PEFT) methods such as LoRA hu2021lora and QLoRA dettmers2023qlora following layer pruning.
First, from the perspective of resource overhead, layer replacement is more adaptable to hardware resource constraints compared to other methods. Fine-tuning the model using the PEFT methods requires storing the model’s weights, activation values, and the optimizer state of the PEFT module in the GPU during training. In contrast, layer replacement involves two stages: dataset construction and model training. The first stage only requires storing the model’s weight and the forward computation overhead, while the second stage requires only the storing of a lightweight network. Therefore, layer replacement can also be implemented under hardware resource constraint conditions.
Second, layer replacement with a lightweight network, which substitutes for pruned layers, ensures that the pruned model more closely resembles the original model compared to models derived using PEFT methods. This similarity leads to a higher potential for performance recovery in the pruned model. This will be validated in our experiments.
@@ 3 Metrics for Evaluating Pruned Models
Accuracy is the most commonly used metric when evaluating LLMs in classification tasks. However, accuracy may overestimate the performance of the model after compression, since it does not take into account the consistency of the model’s answers before and after compression. In this section, we analyze such limitation and propose a novel metric for compressed models.
@@ 3.1 Shortcoming of Accuracy Metric
When evaluating the natural language understanding ability of LLMs, most existing benchmarks formalize the task as a classification task. The input to these tasks is a question accompanied by several choices, and the LLM is required to select the correct answer from these choices. Typically, each choice is concatenated with the question to form multiple sentences, and the perplexity (PPL) of each is computed. The choice corresponding to the sentence with the lowest PPL is selected as the answer. A classification task with $k$ choices and consisting of $N$ samples is denoted as $\mathcal{T}=\{\left(x_{i},c_{i,1},c_{i,2},...,c_{i,k},y_{i}\right)\}_{i=1}^{N}$, where $x_{i}$ represents the question in the $i$-th sample, $c_{i,j}$ represents the $j$-th choices, and $y_{i}$ represents the correct choice. Additionally, we define $\mathcal{M}$ to denote the original LLM, $\bar{\mathcal{M}}$ to denote the compressed LLM, and $\hat{y}(\mathcal{M})$ to denote the choice predicted by the model $\mathcal{M}$. We initially evaluate the performance of the model pruned by the method described in Sec 2.2 and observe the performance of the pruned model has been improved on some classification tasks. To further investigate this phenomenon, we analyze the experimental results using the confusion matrix li2024evaluating . Specifically, we count the number of samples and average standard deviation (std) for the PPL of the samples for each term of the confusion matrix. The calculation of the std for the PPL of the $i$-th sample is defined as follows:
$\small\mathrm{PPL}_{i,j}=\mathrm{PPL}(\mathcal{M}(x_{i},c_{i,j})),{\mathrm{PPL}}_{i}=\frac{\sum_{j=1}^{k}\mathrm{PPL}_{i,j}}{k},\mathrm{std}_{i}=\sqrt{\frac{\sum_{j=1}^{k}(\mathrm{PPL}_{i,j}-{\mathrm{PPL}}_{i})^{2}}{k-1}},$ (5)
@@
Table 1 presents the counts of samples in TP, FN, FP, TN, and also the PPL standard deviation in several datasets. We can observe that samples in ${\rm FP}$ constitute a considerable proportion of the total samples, and the std for ${\rm TP}$ and ${\rm TN}$ is significantly higher than that for ${\rm FN}$ and ${\rm FP}$. This indicates that the model is more uncertain about the ${\rm FN}$ and ${\rm FP}$ samples, implying that the model may guess the correct answer for a significant portion of these samples after pruning. This phenomenon suggests that the accuracy metric may overestimate the performance of the compressed model.

### v4

@@ Abstract
This paper introduces LLM-Streamline, a pioneer work on layer pruning for large language models (LLMs). It is based on the observation that different layers have varying impacts on hidden states, enabling the identification of less important layers to be pruned. LLM-Streamline comprises two parts: layer pruning, which removes consecutive layers with the lowest importance based on target sparsity, and layer replacement, a novel module that trains a lightweight network to replace the pruned layers to mitigate performance loss. Additionally, a new metric called stability is proposed to address the limitations of the widely used accuracy metric in evaluating model compression. Experiments show that LLM-Streamline outperforms both previous and concurrent state-of-the-art pruning methods in terms of both performance and training efficiency.
Figure 1: The left side of the figure illustrates the LLM-Streamline workflow, which includes layer pruning to remove consecutive layers and layer replacement where a lightweight network is trained to replace the pruned layers. The right side of the figure presents the comparison results of LLM-Streamline with the state-of-the-art (SOTA) pruning methods on 12 classification benchmarks (details in Section 4.2) after pruning about 25% of the parameters on Llama2-7B. LLM-Streamline achieves 11.2% higher relative accuracy than these methods, where the relative accuracy represents the percentage of the original model’s accuracy retained by the pruning method.
@@ 1 Introduction
Large language models (LLMs) built on the Transformer architecture (Vaswani et al., 2017) have gained widespread attention and are applied across diverse domains and tasks. However, as LLMs increase in size, their hardware requirements escalate substantially, thereby constraining their applicability and impeding their deployment in real-world scenarios. To reduce the hardware requirements for deploying LLMs, research efforts have been devoted to developing compact models that maintain high performance through model compression (Zhu et al., 2023; Wang et al., 2024). Currently, model compression techniques are widely categorized into knowledge distillation (Hinton et al., 2015; Gou et al., 2021; Li et al., 2022; Huang et al., 2022; Ho et al., 2022), quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022), and pruning (Louizos et al., 2017; Chen et al., 2023; Frantar & Alistarh, 2023; Das et al., 2023; Sun et al., 2023; Xia et al., 2023). Knowledge distillation achieves compression by transferring the capabilities of a larger teacher model to a smaller student model. Quantization compresses the model by quantizing the weights to lower precision. Alternatively, pruning compresses the model by eliminating unimportant parameters and modules.
@@
This work focuses on the popular pruning methods. Previous approaches for LLM primarily prune dense matrices (Ashkboos et al., 2024), attention heads (Michel et al., 2019; Voita et al., 2019), filters (McCarley et al., 2019; Prasanna et al., 2020), or prune parameters to reduce an LLM’s hidden dimension (Xia et al., 2023; van der Ouderaa et al., 2023; Hu et al., 2024). Despite their effectiveness, these methods often result in structural irregularities, making it inflexible to store and deploy the pruned models. In contrast, layer pruning method simply reduces the depth of LLMs. As the layers of LLMs are stored in data structures like nn.ModuleList, layer pruning only requires removing elements from this list, making it more flexible for application. Therefore, exploring an effective layer-wise pruning method is crucial.
Figure 2: The cosine similarity between the input and output hidden states of each layer in OPT-1.3B, OPT-2.7B, OPT-6.7B, and Llama2-7B.
The core idea of layer pruning is to identify and remove less important layers in an LLM. Specifically, the effect of each layer can be viewed as a transformation of the hidden states. If the input and output hidden states of a particular layer are highly similar, such as exhibiting high cosine similarity, we can say that the layer has a small impact on adjusting the hidden states. As illustrated in Fig. 2, our pilot study shows that certain contiguous layers indeed have smaller impact on the hidden states, indicating they are less important and suitable for pruning. Some concurrent works (Song et al., 2024; Kim et al., 2024; Yang et al., 2024; Men et al., 2024; Gromov et al., 2024) also explore layer pruning. These studies either prune unimportant layers directly without further training (Song et al., 2024; Men et al., 2024) or fine-tune the pruned model to enhance performance (Kim et al., 2024; Yang et al., 2024; Gromov et al., 2024). Directly removing layers can lead to more performance degradation. While parameter-efficient fine-tuning techniques like LoRA (Hu et al., 2021) are used to train the pruned LLM, fine-tuning the model to make the original non-contiguous layers compensate for the performance degradation is not an easy task (details in Section 2.3).
In this work, we propose a layer pruning method called LLM-Streamline, which exhibits advantages in both performance and training efficiency while requiring less training data. LLM-Streamline comprises two components: layer pruning and layer replacement. According to a certain target sparsity, the first step removes consecutive layers with the lowest importance from the original LLM. Subsequently, we train a lightweight network to replace the pruned layers, aiming to recover the performance degradation caused by pruning. We can employ various architectures for this lightweight network, including a feed-forward neural network (FFN), a SwiGLU-based feed-forward neural network (SwiGLU), and a Transformer layer.
@@
Additionally, we find that existing accuracy metrics for evaluating model compression methods have limitations. Specifically, in natural language understanding tasks that involve multiple-choice classification, a compressed model may guess correct answers for samples on which the original model was uncertain, resulting in an overestimation of the compression performance. To address this issue, we propose a new metric named stability, which measures the consistency of predictions before and after pruning, considering the prediction confidence of the original model.
@@
  • We propose LLM-Streamline, a layer-wise pruning algorithm that demonstrates superior effectiveness and efficiency compared to concurrent methods. To mitigate the potential performance degradation caused by pruning, we propose to use a lightweight network to approximate the functionality of the pruned layers.
  • We propose a new metric called stability, which considers both the prediction confidence of the original model and the consistency of predictions before and after pruning. Stability provides a more accurate reflection of the pruned model’s performance in classification tasks compared to the widely used accuracy metric.
  • We conduct experiments on 12 well-known classification benchmarks and 3 generation benchmarks. Our results show that for an LLM with 7B or 13B parameters and a 25% pruning rate, we can maintain 93% performance in classification tasks and 77% in generation tasks without requiring a lot of training data, outperforming existing SOTA pruning methods.
@@ 2 LLM-Streamline
The workflow of the LLM-Streamline framework, shown in Fig. 1 (a), comprises two main steps: layer pruning and layer replacement. First, we prune redundant layers from the LLMs. Then, we train a lightweight network to replace the pruned layers to restore the model’s performance.
@@ 2.1 Layer Redundancy in LLMs
LLMs primarily utilize a Transformer architecture, consisting of a series of Transformer decoder layers. These layers adopt a residual structure, so the effect of each Transformer layer can be viewed as a transformation of the input hidden states. Assuming that the parameters of the $\ell$-th layer $f$ are denoted as $\theta^{(\ell)}$, and its input hidden states are represented by $\bm{x}^{(\ell)}$, the layer $f$ can be expressed as
$\bm{x}^{(\ell+1)}=\bm{x}^{(\ell)}+f(\bm{x}^{(\ell)},\theta^{(\ell)}).$ (1)
@@
In Eq. 1, the $\ell$-th layer $f$ contributes a transformation $f(\bm{x}^{(\ell)},\theta^{(\ell)})$ to the input $\bm{x}^{(\ell)}$. Therefore, we assess the importance of each layer in LLMs by evaluating its impact on the input hidden states. We use the cosine similarity $\cos{(\cdot,\cdot)}$ between input $\bm{x}^{(\ell)}$ and output $\bm{x}^{(\ell+1)}$ as a metric. Essentially, a higher cosine similarity between the input and output of a layer indicates lower importance, and vice versa. This interpretation arises from the observation that a high cosine similarity suggests the layer’s transformation is minimal, making it more amenable to pruning.
@@
To measure the importance of different layers in LLMs, we randomly select samples from the pre-training data (details in Section 4.1). We then record the hidden states generated by the LLMs for these samples and compute the cosine similarity between the input and output hidden states of each layer. The computation of cosine similarity can be formalized as follows,
$\cos{(\bm{x}^{(\ell)},\bm{x}^{(\ell+1)})}=\mathbb{E}_{(\bm{x}^{(\ell)}_{i},\bm{x}^{(\ell+1)}_{i})\in\mathcal{D}}\left(\frac{1}{L}\sum_{j=1}^{L}\frac{\bm{x}^{(\ell)}_{i,j}\cdot\bm{x}^{(\ell+1)}_{i,j}}{\|\bm{x}^{(\ell)}_{i,j}\|\cdot\|\bm{x}^{(\ell+1)}_{i,j}\|}\right),$ (2)
@@
where $\mathcal{D}$ denotes the recorded hidden states from different samples, $\bm{x}^{(\ell)}_{i}$, $\bm{x}^{(\ell+1)}_{i}\in R^{d\times L}$ denotes the input and output hidden states of the $i$-th sample respectively, $d$ denotes the hidden size and $L$ denotes the sequence length of each sample.
@@
To mitigate the effects of model size and model structure, we conduct experiments on four models OPT-1.3B, OPT-2.7B, OPT-6.7B (Zhang et al., 2022), and Llama2-7B (Touvron et al., 2023). The results, illustrated in Fig. 2, show that there is high cosine similarity between the input and output of several consecutive layers in all models, indicating a low level of importance.
Discussion I: Why not use other similarity to measure the importance of layers? In deep learning, cosine similarity is widely employed to measure the similarity between two vectors (Chen et al., 2020; Chen & He, 2021; Reimers, 2019). Alongside it, dot product and Euclidean distance are also utilized, but they additionally consider vector magnitude. Current research suggests that the hidden states of Transformers with the Pre-Norm architecture tend to grow as the depth of layers increases (Liu et al., 2023). This trend leads to a bias where deeper layers in the model have higher dot product similarity, while earlier layers have smaller Euclidean distances. Consequently, we opt for cosine similarity, which is agnostic to the magnitude of the vectors.
Discussion II: Why not use perplexity as the metric to measure the importance of layers? Some concurrent layer pruning work uses perplexity as the metric to measure the importance of layers (Song et al., 2024; Kim et al., 2024). Specifically, they remove each layer one at a time, measuring the change in perplexity of the model on the pre-training data, and eliminate the layer that causes the least change. This process is repeated multiple times to remove several layers. However, we think perplexity is a highly data-sensitive metric, which results in different layers being removed when pruning with different pre-training data. This also results in a situation where, although the perplexity of the pruned model on the pre-training data used for pruning is low, it performs poorly on other datasets. In contrast, the cosine similarity is highly stable and always leads to the same pruned layers on different pre-training data. We conduct detailed experiments in the Appendix A to demonstrate that perplexity is a highly data-sensitive metric and performs poorly on downstream tasks.
@@ 2.2 Layer Pruning
As Fig. 2 shows, the less important layers are often contiguous. Hence, given number of pruned layers $n$ determined by a target sparsity, we remove $n$ contiguous layers by finding the initial layer $\ell^{*}(n)$ corresponding to the highest cosine similarity for pruning:
$\ell^{*}(n)=\underset{\ell}{\arg\max}\,\cos(\bm{x}^{({\ell})},\bm{x}^{({\ell+n})}),$ (3)
where we randomly select samples from the pre-training data to compute the cosine similarity between $\bm{x}^{(\ell)}$ and $\bm{x}^{(\ell+n)}$, as outlined in Eq. 2.
@@ 2.3 Layer Replacement
After the layer pruning process, we aim to replace the pruned layer with a lightweight network that has much fewer parameters. The rationale is that these layers contribute only minor transformations to the input. Therefore, we hypothesize that the cumulative effect of these layers can be approximated by a lightweight network. Specifically, after identifying the initial layer $\ell^{*}(n)$ for pruning, we use $(\bm{x}^{(\ell^{*})},\bm{x}^{(\ell^{*}+n)})$ as the training data to train the lightweight network using mean squared error (MSE) loss, which can be formalized as follows:
@@
Discussion: Layer Replacement or Fine-Tuning Pruned LLMs? Here, we discuss why opt for layer replacement, instead of using common Parameter-Efficient Fine-Tuning (PEFT) methods such as LoRA  (Hu et al., 2021) and QLoRA  (Dettmers et al., 2023) after layer pruning.
First, from the perspective of resource overhead, layer replacement is more adaptable to hardware resource constraints compared to other methods. Fine-tuning the model using the PEFT methods requires storing the model’s weights, activation values, and the optimizer state of the PEFT module in the GPU during training. In contrast, layer replacement involves two stages: dataset construction and model training. The first stage only requires storing the model’s weight and the forward computation overhead, and the second stage only requires storing of the lightweight network’s weight, activation values of lightweight network, and the optimizer state of lightweight network. Therefore, layer replacement can also be implemented under conditions of hardware resource constraints.
Second, layer replacement uses a lightweight network to replace the pruned layer, and distills the knowledge of the pruned layer into the lightweight network using the MSE loss function. Unlike layer replacement, we speculate that training the model after pruning with LoRA is a process of redistributing the function of the pruned layers across the remaining layers. Therefore, substituting the pruned layers with a lightweight network could be a less challenging training task than redistributing the function of the pruned layers across the remaining layers. In the experiments of Section 4.7, we demonstrate that layer replacement has better performance compared to LoRA.
@@ 3 Metrics for Evaluating Pruned Models
Accuracy is the most commonly used metric for evaluating LLMs in classification tasks. However, accuracy may overestimate the performance of the model after compression, since it does not take into account the consistency of the model’s answers before and after compression. In this section, we analyze such limitation and propose a novel metric for evaluating compressed models.
@@ 3.1 Shortcoming of Accuracy Metric
When evaluating the natural language understanding capabilities of LLMs, most existing benchmarks frame the task as a classification task. A classification task with $k$ choices and comprising $N$ samples is denoted as $\mathcal{T}=\{\left(x_{i},c_{i,1},c_{i,2},...,c_{i,k},y_{i}\right)\}_{i=1}^{N}$, where $x_{i}$ represents the question in the $i$-th sample, $c_{i,j}$ represents the $j$-th choices, and $y_{i}$ represents the correct choice. The input to the classification task consists of a question accompanied by multiple choices, and the LLM is required to select the correct answer from these choices. Typically, each choice is concatenated with the question to form multiple sentences, and the perplexity (PPL) of each sentence is computed. The choice corresponding to the sentence with the lowest PPL is selected as the answer.
Typically, model pruning results in decreased model performance. However, when we evaluate the model pruned by the method described in Sec 2.2, we unexpectedly observe the accuracy of the pruned model has been improved on some classification tasks. We define $\mathcal{M}$ to denote the original LLM, $\bar{\mathcal{M}}$ to denote the compressed LLM, and $\hat{y}(\mathcal{M})$ to denote the choice predicted by the model $\mathcal{M}$. To further investigate this phenomenon, we analyze the experimental results using the confusion matrix  (Li et al., 2024). Specifically, we count the number of samples and average standard deviation (std) for the PPL of the samples for each term of the confusion matrix. The calculation of the std for the PPL of the $i$-th sample is defined as follows:
$\mathrm{PPL}_{i,j}=\mathrm{PPL}(\mathcal{M}(x_{i},c_{i,j})),{\mathrm{PPL}}_{i}=\frac{\sum_{j=1}^{k}\mathrm{PPL}_{i,j}}{k},\mathrm{std}_{i}=\sqrt{\frac{\sum_{j=1}^{k}(\mathrm{PPL}_{i,j}-{\mathrm{PPL}}_{i})^{2}}{k-1}},$ (5)
@@
Table 1 presents the counts of samples in TP, FN, FP, TN in sevaral datasets, and also the PPL standard deviation in Race-H dataset. We can observe that the std for ${\rm TP}$ and ${\rm TN}$ is significantly higher than that for ${\rm FN}$ and ${\rm FP}$. This indicates that the model is more uncertain about the ${\rm FN}$ and ${\rm FP}$ samples. In addition, the samples in ${\rm FP}$ constitute a considerable proportion of the total samples, implying that the model may guess the correct answer for a significant portion after pruning. This phenomenon suggests that the accuracy metric may overestimate the performance of the compressed model.